OpenStudy (anonymous):

which value is a solution to the equation 2-8x=-6? 1 1/2 -1 8

OpenStudy (carolina4567):

A = 1 Thats ur answer

OpenStudy (anonymous):

I know i was just using a trick question :) hoping somebody would know

OpenStudy (carolina4567):

2 - 8 x = -6 +2 +2 8x = 8 8 8 x = 1